The activity at the Soufriere Hills Volcano during this period has been at a slightly higher level than that observed during the previous 24 hours.
Seismicity continues to be dominated by small- to moderate-sized rockfall signals. None of these was very large, although light ashfall was generated on a number of occasions. The largest of these events occurred at about 14:00 on 25 May and produced ashfall in Corkhill and Old Towne. Ash clouds generated during the period were thinner than has recently been the case, and showed very little thermal convection. The pattern of events experienced overnight and today is quite similar to earlier periods of phreatic activity at the Soufriere Hills Volcano and suggests that heavy rainfall between 03:00 and 06:00 on 25 May, may have been caused some increase in phreatic activity.
There were 17 long-period events and one VT earthquake recorded by the seismic network today. The VT earthquake was located at a depth of about 2.5 km beneath English's Crater. The level of broadband tremor increased during this reporting period, and there was a sustained period of low- to medium-amplitude broadband tremor between 02:59 and 06:17 on 25 May. The tremor became more intermittent after that. A number of long duration seismic signals possibly due to mudflows were recorded last night and today. The signals are much longer than typical rockfall signals, lasting for up to 20 minutes. They show much greater variation in amplitude than the broadband tremor.
Visibility was very poor, and the dome was covered in cloud most of the day. There was a brief period around lunchtime when partial views of the dome were possible from a number of locations. There are now at least three spines present on the dome, although none of them are more than 50 ft high. Vigorous steaming was seen in the northwest moat, and from several other areas on the dome. A helicopter inspection of the recent pyroclastic flow deposits revealed the presence of a new deposit which had descended the Upper Tar River Valley down to approximately 1500 ft from the base of the dome. The deposit appear to consist of remobilised wet ash, and may be a mudflow caused by the heavy rainfall last night. The lower eastern flank of the dome was also visible during the inspection. There was a clear scar, low on the eastern flank, about 3 ft deep and perhaps 15 to 30 ft wide. A number of rockfalls were seen travelling down this feature.
No EDM or GPS measurements were made today, due to poor visibility and equipment failure respectively.
Further COSPEC measurements were made today and the data is still being processed. No FTIR measurements were made. The FTIR measurements made on 24 May in Plymouth Docks and Fox's Bay indicate that HCL and SO2 levels are still well below levels of concern for human habitation.
Dr Jenni Barclay from Bristol University left the island this afternoon after a five-week visit. She has been working on mapping the dome and estimating its volume. Dr Pierre Delmelle from Brussels University also left; he spent his time here making COSPEC measurements and sampling fluids at the various soufrieres. Dr Simon Young of the BGS arrives on Montserrat this afternoon for a one month tour of duty. Dr Rosalyn Lopez, Jet Propulsion Laboratories, Dr Haruldur Sigurdsson, University of Rhode Island, Professor Wagner, University of Geneva, and Dr Laurent Stieltjes, BRGM, France, all arrived today for short fact finding visits to MVO.
The present level of activity at the Soufriere Hills Volcano continues to cause concern to the scientists. The MVO urges that visits to the evacuated zone are kept to a minimum. The Tar River, Long Ground and Whites areas are extremely dangerous and should not be entered under any circumstances.